Please copy and paste your Excel bond calculations under each relevant question. Do NOT submit an Excel spreadsheet, only your word document with your name on it.

A) A 30-year Treasury bond expiring on February 15, 2048 with a 3% coupon has a yield of 5%. Calculate its price. Assume the bond’s settlement date is June 30th, 2018.

B) You want to buy a bond that pays an annual coupon of 4.2% on March 31st of each year. On June 30th, 2018, the bond is priced at 99-12 (in points and fractions). You decide to buy (and pay for) the bond on that date at that price. What is the final invoice price that you must pay to the bond seller?

C) Consider an original 5-year US Treasury note with a 3.5% coupon expiring on 06/15/2022. If this note sells at 101-07 in the bond market, what would be its yield to maturity? Assume the bond’s settlement date is 6/30/2018. Remember that US government bonds and notes pay the coupon on a semi-annual basis)

D) General Motors recently issued a 7-year note that pays a 4% coupon rate (paid annually) and expires on February 15, 2025. The current price of this note is 100-04 (in points and fractions). What is this note’s yield to maturity? And assume that the bond’s settlement date is 6/30/2018.

E) It is now one year later and market yields for bonds/notes with similar characteristics have risen to 4.5%. What should now be the (dollar) price of the Albertson’s note? How do you explain the change of price from one year to the other?
Answer #1

Soln : 1) Please refer the table :

Year 1 2.0000 3 4.0000 5 6.0000 7 8.0000 9 10.0000 11 12.0000 13 14.0000 15 16.0000 17 18.0000 19 20.0000 21 22.0000 23 24.0000 25 26.0000 27 28.0000 29 30.0000
Cash flow 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 30.00 1030.00
Discount at 5% 0.95 0.91 0.86 0.82 0.78 0.75 0.71 0.68 0.64 0.61 0.58 0.56 0.53 0.51 0.48 0.46 0.44 0.42 0.40 0.38 0.36 0.34 0.33 0.31 0.30 0.28 0.27 0.26 0.243 0.236
PV 28.57 27.21 25.92 24.68 23.51 22.39 21.32 20.31 19.34 18.42 17.54 16.71 15.91 15.15 14.43 13.74 13.09 12.47 11.87 11.31 10.77 10.26 9.77 9.30 8.86 8.44 8.04 7.65 7.29 242.66
Sum 696.89

Price of the bond = $ 696.89

B) Bond is priced at 99-12 i.e. 99*1000/100 +12/32*1000/100 = 993.75. As bond is to be bought on 30th June 2018

Coupon = 4.2% , Here the buyer will pay the accrued interest also on this bond

Accrued interest = 42 *270/360 = 31.5

Invoice value that seller will get = 993.75 + 31.5 = $1025.25

3) Let r be the yield to maturity, Price of the bond = 101*10 + 2.1875 = 1012.1875

Accrued interest = 17.5*168/360 = 8.16

Actual price = 1012.1875-8.16 = 1004

On solving by hit & trial we get r = 1.71*2 = 3.42%

d) Similarly here we will calculate the yield rate R , price = 1000 + 1.25 = 1001.25

We will get the yield rate as R = 4% (approx.)

Year 1 2.0000 3 4.0000 5 6.0000 7 8.0000 9 10.0000 11 12.0000 13 14.0000
Cash flow 40.00 40.00 40.00 40.00 40.00 40.00 40.00 40.00 40.00 40.00 40.00 40.00 40.00 1040.00
Discount factor 0.96 0.92 0.89 0.85 0.82 0.79 0.76 0.73 0.70 0.68 0.65 0.62 0.60 0.58
PV 38.46 36.98 35.56 34.19 32.88 31.61 30.40 29.23 28.10 27.02 25.98 24.98 24.02 601.52
Sum 1000.94